(2x-3)/(4)-(x-5)/(6)=1
Answer by stanbon(75887)

(2x-3)/(4)-(x-5)/(6)=1
---
Multiply thru by 12
---
3(2x-3) - 2(x-5) = 12
---
6x-9-2x+10 = 12
4x+1 = 12
4x = 11
x = 11/4
=======================
Cheers,
Stan H.